OpenStudy (anonymous):

Solve for x: |x - 2| + 10 = 12

OpenStudy (anonymous):

x-2 = 0 x =2 now you take case where x < 2 and x >= 2 x<2 so : -(x-2) + 10 = 12 -x +2 +10 = 12 -x = 0 x = 0 x >= 2 : (x-2) + 10 = 12 x +8 = 12 x = 4 so x = 4 , 0